Vector A=2ti+tj-t^3k and B=sinti+costj evaluate

A..d/dt(A.B)

B..d/dt(A.A)

C..d/dt(A×B)

D..show that d/dt(A×A) is equal to zero.


1
Expert's answer
2022-02-14T18:33:30-0500

A.


"A\cdot B=2t\sin t+t\cos t"

"\dfrac{d}{dt}(A\cdot B)=2\sin t+2t\cos t+\cos t-t\sin t"

B.



"A\cdot A=4t^2+t^2=5t^2"

"\dfrac{d}{dt}(A\cdot A)=10t"

C.


"A\times B=\begin{vmatrix}
 i & j & k \\
 2t & t & -t^3 \\
 \sin t & \cos t & 0 \\
\end{vmatrix}"

"=i(0+t^3\cos t)-j(0+t^3\sin t)+k(2t\cos t-t\sin t)"

"\dfrac{d}{dt}(A\times B)=(3t^2\cos t-t^3\sin t)i-(3t^2\sin t+t^3\cos t)j"

"+(2\cos t-2t\sin t-\sin t-t\cos t)k"

D.


"A\times A=\begin{vmatrix}
 i & j & k \\
 2t & t & -t^3 \\
 2t & t & -t^3 \\
\end{vmatrix}"

"=i(-t^4+t^4)-j(-2t^4+2t^4)+k(2t^2-2t^2)=0"

"\dfrac{d}{dt}(A\times A)=0"